OpenStudy (anonymous):

consider the differential equation dy/dx=2x-y. Find (dy)^(2)/(dx)^(2) in terms of x and y. 

OpenStudy (turingtest):

\[\frac{(dy)^2}{(dx)^2}=(\frac{dy}{dx})^2\]???not sure what you mean...

OpenStudy (anonymous):

do you mean the second derivtive?